OpenStudy (anonymous):

divide by long division (x^2+9x+12)/x+7

OpenStudy (psymon):

It's a process of divide, multiply, subtract, divide, multiply, subtract. It's really those same 3 steps over and over again, it just needs to be done properly.

OpenStudy (anonymous):

I got x+2+(-2/x+7)

OpenStudy (psymon):

So this is what we have: x+7|x^2 + 9x + 12 So we are going to start with the first term in the quotient, x^2. What we do is divide x^2 by the x we have. So that is done like this: (x^2/x) = x. This we put on top: x x+7|x^2 + 9x + 12 So as I mentioned, this is 3 steps, divide, multiply, subtract. That was the divide part. Next is multiply. So the x we have on top. This must be multiply by the whole thing we are dividing by, mainly (x+7). When we do that multiplication, we end up with x(x+7) = x^2 + 7x. This we put underneath like this: x x+7|x^2 + 9x + 12 x^2 + 7x Now this is the subtraction part. We must subtract what we just put underneath from what's above it. So doing that I get x x+7|x^2 + 9x + 12 -(x^2 + 7x) = 2x + 12 So the x^2's cancel, i subtracted 7x from 9x to get 2x and I brought down the 12. Ill stop here for now, can you make sense of what Ive been doing?

OpenStudy (anonymous):

Yes thank you.

OpenStudy (psymon):

Do you think you could repeat the pricess with the 2x + 12?

OpenStudy (anonymous):

yes